Two charges 2.0x10^(-6) C and -1.0 x 10^(-6) C are placed at a separation of 10 cm . Locate the position where a third charge will not experience a net force.

Grade:

1 Answers

Eshan
5 years ago
A third charge won’t experience a force where the net electric field due to the two charges is zero.

Let the point be at a distancexfrom the negative charge along the line joining the charges and away from the direction of positive charge.

Hence at that point, the electric fields due to two charges should cancel each other.

\implies \dfrac{k(2\times 10^{-6})}{(10+x)^2}+\dfrac{k(-1\times 10^{-6})}{x^2}=0
\implies x=24.14cm